OpenStudy (anonymous):

Which of the following expressions are equal to the one below? Check all that apply. 5 (8 - 2) A. 5 8 - 5 2 B. (5 8) - 2 C. (8 - 2) 5 D. 8 5 - 2

OpenStudy (carniel):

@Hoblos go for it :O

OpenStudy (hoblos):

5*(8 - 2) = 5*8 - 5*2 = (8-2)*5

OpenStudy (anonymous):

A. C

OpenStudy (hoblos):

these are destributive and cummutative properties